I need to use quadratic equations to find two real numbers that satisfy the sum of 15 and the product 36
Found 2 solutions by mananth, amoresroy:
Answer by mananth(16949)

x+y =15
xy = 36
x= 36/y
..
plug value of x in the first equation
36/y +y =15
(36+y^2)/y =15
36+y^2 =15y
y^2-15y+36=0
...
y^2-12x-3x+36=0
y(y-12)-3(y-12)=0
(y-12)(y-3)=0
y =12 OR y =3
...
m.ananth@hotmail.ca

Answer by amoresroy(361)

I need to use quadratic equations to find two real numbers that satisfy the sum of 15 and the product 36
Let x = the 1st number
y = the 2nd number
x+y = 15
xy = 36
y = 36/x
x+36/x = 15
multiply both sides by x
x^2+36 = 15x
x^2-15x+36 = 0
(x-3)(x-12) = 0
x = 3
y = 12
The numbers are 3 and 12